The bank charges a nominal interest rate is 16,5% per annum, compounded at the end of every second month. The equivalent effective interest rate is


[1] 17,18%.

[2] 17,68%

. [3] 14,53%

. [4] 16,18%.

[5] none of the above. 


1
Expert's answer
2021-05-24T19:17:09-0400

"1+i=(1+r\/m)^m"

i=effective interest rate, m=6 compound periods

"1+i=(1+0.165\/6)^6=1.1768"

"i=1.1768-1=0.1768=0.1768\times100\%=17.68\%"